Solve with the slope-intercept form of equation (-7,7) and(0,-3)
Answer by mananth(12270)

Solve with the slope-intercept form of equation (-7,7) and(0,-3)
...
slope = m = (-3-(7))/ ((0-(-7))
m=-10/7
...
find b in y= mx +b by plugging (0,-3) & m
...
y=mx+b
-3=0+b
b=-3
...
The equation will be
y=(-10x/7)-3
...
m.ananth@hotmail.ca